# Alexandrov spaces are CS sets

### v2

@@ Abstract.
We prove that the extremal stratification of an Alexandrov space introduced by Perelman–Petrunin is a CS stratification in the sense of Siebenmann. We also show that every space of directions of an Alexandrov space without proper extremal subsets is homeomorphic to a sphere. In the polyhedral case, we obtain a PL structure.
@@ 1.1. Main result
In [29, p.6], [30, p.207], Perelman only stated that the MCS stratification of an Alexandrov space is a WCS stratification (Arrow (4)), which was enough for him to develop the Morse theory with the help of the deformation theorem of Siebenmann for WCS sets. However, there was no mention of the CS structure in [29], [30], even in [33], and as already remarked, since then there have been no papers about it (except for a closely related result of Kapovitch [23, §9] mentioned below). This is why we consider this problem, and the answer is, as stated in Theorem 1.1, given by the extremal stratification of Perelman–Petrunin (Arrow (1)).
@@ 1.2. Further results
Now we discuss further regularity results on the extremal stratification of Alexandrov spaces.
@@
We shall verify Conjecture 1.3 in the case of polyhedral spaces. In this case, the conclusion of Conjecture 1.3 is equivalent to the condition that the associated triangulation is PL. See Section 5 for more details.
@@
If a polyhedral Alexandrov space contains no proper extremal subsets, then the triangulation is PL.
@@ 2.2. Alexandrov spaces
Fix $\kappa\in\mathbb{R}$. The model plane of curvature $\kappa$ refers to a sphere, Euclidean plane, or a hyperbolic plane of constant curvature $\kappa$. For a geodesic triangle $\triangle pqr$ in a metric space, the comparison triangle ~$\tilde{\triangle}pqr$ means a geodesic triangle in the model plane with the same side-lengths (if it exists and is unique up to isometry when $\kappa>0$).
@@
A complete geodesic space $X$ is called an Alexandrov space with curvature $\geq\kappa$ if for any geodesic triangle $\triangle pqr$ in $X$ and the comparison triangle ~$\tilde{\triangle}pqr$ in the model plane of curvature $\kappa$, the natural correspondence
@@ 2.3. MCS spaces
A metrizable space $X$ is called an MCS space of dimension $n$ if, for any $x\in X$, there exist a neighborhood $U$ of $x$ in $X$ and a compact MCS space $L$ of dimension $n-1$ together with a pointed homeomorphism
@@
where $o$ is the vertex of the cone $C(L)$. Here we assume that the empty set is the unique compact $(-1)$-dimensional MCS space, the cone over which is a point.
@@ 2.6. Double suspension
$F:=\left\{x\in E\mid\text{no open neighborhood in $X$ is isomorphic to $B^{m}\times C(L)$}\right\},$
@@
[…]
@@
Since $r$ is small enough, it follows from the stability theorem that $\partial B(p,r)$ is homeomorphic to $\Sigma_{p}$. In particular, $\partial B(p,r)$ is a homology manifold with isolated nonmanifold points (Property (2)). Since $\varepsilon$ is small enough, it follows from the fibration theorem (with a gluing argument) that $\partial B(p,r)\cap B(x,\varepsilon)$ is homeomorphic to the open cone $C(\Sigma)$. Thus $\Sigma$ is simply-connected if and only if $x$ is a manifold point of $\partial B(p,r)$ (the “only if” part follows from [8], [4]; see also [26, 6.2] and note that $\dim\partial B(p,r)\geq 4$). In summary,
@@
More generally, Theorem 4.1 holds for any Alexandrov homology manifold $X$: the conclusion in this setting is that $E$ is an extremal subset of dimension $\leq 1$ (in particular, $E$ contains the isolated nonmanifold points of $X$). Let us trace the above proof, checking for required modifications. The dimension $\leq 4$ case follows from Properties (1)–(4), so we may assume dimension $\geq 5$ as before. Step 1 of the proof of Theorem 4.1 works exactly the same way. Step 2 requires the following minor modifications. In the proof of Claim 4.3, we used the assumption that $X$ is a topological manifold to ensure that $\Sigma_{x}$ is simply-connected (Property (3)), which implied that $\Sigma_{x}$ is a topological sphere if and only if it is a topological manifold. However, in the proof of Claim 4.3, $\Sigma_{x}$ is a topological suspension over $\Sigma$, where $\Sigma$ is connected since $\partial B(p,r)$ is a homology manifold. Therefore $\Sigma_{x}$ is automatically simply-connected. Another difference is that, in the last paragraph, Claim 4.3 may not imply that $E$ has no isolated points. We have to check that if $E$ has an isolated point $p$, then it is an extremal point. Indeed, Claim 4.3 implies that $\Sigma_{p}$ is a topological manifold, and thus it is not simply-connected (otherwise, it is a topological sphere). By Property (3), $p$ is an extremal point.
@@
[…]
@@
Let $\Sigma$ be an Alexandrov homology manifold with curvature $\geq 1$. If $\Sigma$ contains four points with pairwise distance $>\pi/2$, then every space of directions is homeomorphic to a sphere.
Here $\Sigma$ itself is also a topological sphere, in particular, a topological manifold; see Remark 4.10(2). Hence Corollary 4.6 is an immediate consequence of Theorem 4.1 and the $k=1$ case of the following general lemma. By convention, the empty set has dimension $-1$.
@@
Let $\Sigma$ be an Alexandrov space with curvature $\geq 1$. Suppose $\Sigma$ contains a $k$-dimensional extremal subset $F$, where $-1\leq k\leq\dim\Sigma$, satisfying the additional conditions as in Lemma 4.2 when $k=-1,0$. Then $\Sigma$ does not contain $k+3$ points with pairwise distance $>\pi/2$.
@@
Let $C(\Sigma)$ be the Euclidean cone over $\Sigma$. Then the subcone $C(F)$ is a $(k+1)$-dimensional extremal subset (this follows, for example, Lemma 4.2). If $\Sigma$ contains $k+3$ points with pairwise distance $>\pi/2$, then there exists a regular map at the vertex of $C(\Sigma)$ to $\mathbb{R}^{k+2}$. This contradicts Remark 2.20. ∎
@@
In general, if an Alexandrov space $\Sigma$ with curvature $\geq 1$ contains $k+1$ points with pairwise distance $>\pi/2$, then $\Sigma$ is homeomorphic to a $k$-fold suspension ([17, Theorem C]). In the maximal case, i.e., $k=\dim\Sigma+1$, we obtain that $\Sigma$ is bi-Lipschitz homeomorphic to the standard sphere (cf. [29, 3.2], [30]). In the Riemannian case, the existence of such $\pi/2$-discrete points gives some restriction on the differentiable structure ([18]).
@@
    $\Sigma$ contains no nonmanifold points and has diameter $>\pi/2$; or
@@
    $\Sigma$ contains three points with pairwise distance $>\pi/2$.
@@ 5. Polyhedral case
A polyhedral space $X$ is PL (piecewise linear) if the link around every simplex $\sigma$ in $X$ is homeomorphic to a sphere. Note that any PL polyhedral space is in particular a topological manifold. If a polyhedral space is PL, then every link and space of directions are again PL. Furthermore, the following holds.
@@
Let $X$ be a polyhedral space. Then $X$ is PL if and only if every iterated space of directions is homeomorphic to a sphere.
Therefore, in the polyhedral case, Conjecture 1.3 is equivalent to Theorem 1.4.
@@
The “if” part follows by taking iterated spaces of directions of both sides of Equation (5.1). The “only if” part is proved by induction on dimension. If $X$ is PL, Equation (5.1) shows that every space of directions is homeomorphic to a sphere. Since the link is also PL, the inductive assumption and Equation (5.1) show that every $k$-th iterated space of directions ($k\geq 2$) is homeomorphic to a sphere. ∎
@@
Let $X$ be a polyhedral Alexandrov homology $n$-manifold. For any integer $1\leq k\leq n$, let $E_{k}$ be the set of points $p\in X$ such that some $k$-th iterated space of directions at $p$ is not homeomorphic to $S^{n-k}$. If $E_{k}$ is nonempty, then it is an extremal subset of $X$ that is a subcomplex of $X$ only with $k$- or $(k-1)$-dimensional maximal simplexes.
@@
The proof only uses the properties of polyhedral spaces and does not rely on the properties of Alexandrov spaces. Let $U$ be a neighborhood of $p$ that is isometric to a neighborhood of the vertex of the $\kappa$-cone over $\Sigma_{p}X$. Let $x\in U$ be a point in the direction $\xi\in\Sigma_{p}X$. Then we have
@@
By the inductive assumption and Claim 5.3, for any $p\in E_{k}(X)$, the space of directions $\Sigma_{p}E_{k}(X)$ is an extremal subset of $\Sigma_{p}X$ that is a subcomplex of $\Sigma_{p}X$ only with $(k-1)$- or $(k-2)$-dimensional maximal simplexes. This implies that $E_{k}(X)$ only have $k$- or $(k-1)$-dimensional maximal simplexes. Furthermore, by Lemma 4.2, $E_{k}(X)$ is an extremal subset of $X$. Note that the additional conditions in Lemma 4.2 are satisfied. Indeed, Claim 5.3 implies that $\Sigma_{p}E_{k}(X)$ is nonempty (since $k\geq 2$), and if $\Sigma_{p}E_{k}(X)$ consists of one element (when $k=2$), the same argument as in Step 1 of the proof of Claim 4.3 shows that $\operatorname{diam}\Sigma_{p}E_{k}(X)\leq\pi/2$. This completes the proof. ∎
@@
Let $\Sigma$ be a polyhedral Alexandrov homology manifold with curvature $\geq 1$. Suppose $\Sigma$ contains $k+3$ points with pairwise distance $>\pi/2$. Then every $k$-th iterated space of directions is homeomorphic to a sphere.
@@
Let $P$ be a truncated cube obtained by cutting off all the corners of a solid cube by planes. Let $p\in P$ a midpoint of an edge of the original cube. Consider the double $D(P)$ of $P$, i.e., the gluing of two copies of $P$ along the boundary. Then $D(P)$ contains no proper extremal subsets. On the other hand, the space of directions $\Sigma_{p}D(P)$ is isometric to the double of the quadrant of the unit sphere, which contains a 00-dimensional extremal subset. A similar example can be obtained by taking the intersection of the solid cube in $\mathbb{R}^{3}$ with edge length $2$ and the concentric Euclidean ball of radius $\sqrt{2}$ and doubling it. In this case the resulting space has no proper extremal subsets even “locally” around the midpoint of the original edge.
@@
The $xw$-, $yw$-, and $zw$-planes are primitive extremal subsets of $X$. In particular, their union $E$ is an extremal subset of $X$, but not primitive. The projection $\pi(E)$ is an extremal subset of codimension $2$ in $X/\mathbb{Z}$ ([33, 4.1]). Furthermore, $\pi(E)$ is primitive because it can no longer be divided into each plane, due to the effect of the group action. However, $\pi(E)$ does not admit any Alexandrov metric with a lower curvature bound as its intrinsic geodesics branch.

### v3

@@ Abstract.
We prove that the extremal stratification of an Alexandrov space introduced by Perelman–Petrunin is a CS stratification in the sense of Siebenmann. We also show that every space of directions of an Alexandrov space without proper extremal subsets is homeomorphic to a sphere. In the polyhedral case, the same holds for every iterated space of directions.
@@ 1.1. Main result
In [29, p.6], [30, p.207], Perelman only stated that the MCS stratification of an Alexandrov space is a WCS stratification (Arrow (4)), which was enough for him to develop the Morse theory with the help of Siebenmann’s deformation theorem for WCS sets. However, there was no mention of the CS structure in [29], [30], even in [33], and as already remarked, since then there have been no papers about it (except for a closely related result of Kapovitch [23, §9] mentioned below). This is why we consider this problem, and the answer is, as stated in Theorem 1.1, given by the extremal stratification of Perelman–Petrunin (Arrow (1)).
@@ 1.2. Further results
Now we discuss further regularity results for the extremal stratification of Alexandrov spaces, which is a new contribution of this paper.
@@
We shall verify Conjecture 1.3 in the polyhedral case. See Section 5 for the definition of a polyhedral space. In this case, the conclusion of Conjecture 1.3 is equivalent to the condition that the link around every simplex is homeomorphic to a sphere (Lemma 5.1).
@@
For polyhedral Alexandrov spaces, Conjecture 1.3 is true.
Remark 1.5.
In the polyhedral case, it would be more natural to expect a PL structure, i.e., the link around every simplex is PL homeomorphic to a sphere.
@@
Remark 1.7.
Although the above theorems only deal with the top stratum of an Alexandrov space, the same results could be extended to each extremal stratum of an Alexandrov space. However, we will not treat this generalization in this paper, as our primary goal is to prove Conjecture 1.3.
@@ 2.2. Alexandrov spaces
Fix $\kappa\in\mathbb{R}$. The model plane of curvature $\kappa$ refers to a sphere, Euclidean plane, or a hyperbolic plane of constant curvature $\kappa$. For a geodesic triangle $\triangle pqr$ in a metric space, the comparison triangle $\tilde{\triangle}pqr$ means a geodesic triangle in the model plane with the same side-lengths (if it exists and is unique up to isometry when $\kappa>0$).
@@
A complete geodesic space $X$ is called an Alexandrov space with curvature $\geq\kappa$ if for any geodesic triangle $\triangle pqr$ in $X$ and the comparison triangle $\tilde{\triangle}pqr$ in the model plane of curvature $\kappa$, the natural correspondence
@@ 2.3. MCS spaces
A metrizable space $X$ is called an MCS space of dimension $n$ if, for any $x\in X$, there exist a neighborhood $U$ of $x$ in $X$ and a compact MCS space $L$ of dimension n1n-1 together with a pointed homeomorphism
@@
where $o$ is the vertex of the cone $C(L)$. Here we assume that the empty set is the unique compact (1)(-1)-dimensional MCS space, the cone over which is a point.
@@ 2.6. Double suspension
[…]
@@
$\iff$ $\text{every space of directions of $\Sigma_{p}$ is simply-connected}.$
@@
Since $r$ is small enough, it follows from the stability theorem that $\partial B(p,r)$ is homeomorphic to $\Sigma_{p}$. In particular, $\partial B(p,r)$ is a homology manifold with isolated nonmanifold points (Property (2)). Since $\varepsilon$ is small enough, it follows from the fibration theorem that $\partial B(p,r)\cap B(x,\varepsilon)$ is homeomorphic to the open cone $C(\Sigma)$ (more precisely, the fibration theorem implies that $\partial B(p,r)\cap\bar{B}(x,t)\setminus B(x,s)$ is homeomorphic to $\Sigma\times[s,t]$ for any $0<s<t<\varepsilon$, and by gluing such homeomorphisms we obtain the desired cone structure). Thus $\Sigma$ is simply-connected if and only if $x$ is a manifold point of $\partial B(p,r)$ (the “only if” part follows from [8], [4]; see also [26, 6.2] and note that $\dim\partial B(p,r)\geq 4$). In summary,
@@
More generally, Theorem 4.1 holds for any Alexandrov homology manifold $X$: the conclusion in this setting is that $E$ is an extremal subset of dimension $\leq 1$ (in particular, $E$ contains the isolated nonmanifold points of $X$). Let us trace the above proof, checking for required modifications.
The dimension $\leq 4$ case follows from Properties (1)–(4), so we may assume dimension $\geq 5$ as before. Step 1 of the proof of Theorem 4.1 works exactly the same way. Step 2 requires the following minor modifications. In the proof of Claim 4.3, we used the assumption that $X$ is a topological manifold to ensure that $\Sigma_{x}$ is simply-connected (Property (3)), which implied that $\Sigma_{x}$ is a topological sphere if and only if it is a topological manifold. However, in the proof of Claim 4.3, $\Sigma_{x}$ is a topological suspension over $\Sigma$, where $\Sigma$ is connected since $\partial B(p,r)$ is a homology manifold. Therefore $\Sigma_{x}$ is automatically simply-connected. Another difference is that, in the last paragraph, Claim 4.3 may not imply that $E$ has no isolated points. We have to check that if $E$ has an isolated point $p$, then it is an extremal point. This is proved as follows. Since $p$ is isolated in $E$, Claim 4.3 implies that $\Sigma_{p}$ is a topological manifold. Thus $\Sigma_{p}$ is not simply-connected (otherwise, it is a topological sphere). By Property (3), $p$ is an extremal point.
@@
$\iff$ $\text{$S(\Sigma_{\xi})$, equivalently $\Sigma_{p_{i}}$, is a topological sphere}.$
@@
Let $\Sigma$ be an Alexandrov homology manifold with curvature $\geq 1$. If $\Sigma$ contains four points with pairwise distance $\pi/2$, then every space of directions is homeomorphic to a sphere.
Here $\Sigma$ itself is also a topological sphere, in particular, a topological manifold; see Remark 4.10(2). Hence Corollary 4.6 is an immediate consequence of Theorem 4.1 and the $k=1$ case of the following general lemma. By convention, the empty set has dimension 1-1.
@@
Let $\Sigma$ be an Alexandrov space with curvature $\geq 1$. Suppose $\Sigma$ contains a $k$-dimensional extremal subset $F$, where $1\leq k\leq\dim\Sigma$, satisfying the additional conditions as in Lemma 4.2 when k=1,0k=-1,0. Then $\Sigma$ does not contain $k+3$ points with pairwise distance $\pi/2$.
@@
Let $C(\Sigma)$ be the Euclidean cone over $\Sigma$. Then the subcone $C(F)$ is a $(k+1)$-dimensional extremal subset (this follows, for example, Lemma 4.2). If $\Sigma$ contains $k+3$ points with pairwise distance $\pi/2$, then there exists a regular map at the vertex of $C(\Sigma)$ to $\mathbb{R}^{k+2}$. This contradicts Remark 2.20. ∎
@@
In general, if an Alexandrov space $\Sigma$ with curvature $\geq 1$ contains $k+1$ points with pairwise distance $\pi/2$, then $\Sigma$ is homeomorphic to a $k$-fold suspension ([17, Theorem C]). In the maximal case, i.e., $k=\dim\Sigma+1$, we obtain that $\Sigma$ is bi-Lipschitz homeomorphic to the standard sphere (cf. [29, 3.2], [30]). In the Riemannian case, the existence of such $\pi/2$-discrete points gives some restriction on the differentiable structure ([18]).
@@
    $\Sigma$ contains no nonmanifold points and has diameter $\pi/2$; or
@@
    $\Sigma$ contains three points with pairwise distance $\pi/2$.
@@ 5. Polyhedral case
We say that a polyhedral space $X$ has spherical links if the link around every simplex in $X$ is homeomorphic to a sphere. If $X$ has spherical links, then $X$ is a topological manifold. Furthermore, the link $L(\sigma)$ of a simplex $\sigma$ in $X$ also has spherical links. The following lemma describes the relation between Conjecture 1.3 and Theorem 1.4.
@@
Let $X$ be a polyhedral space. Then $X$ has spherical links if and only if every iterated space of directions is homeomorphic to a sphere.
@@
The “if” part follows by taking iterated spaces of directions of both sides of Equation (5.1). The “only if” part is proved by induction on dimension. If $X$ has spherical links, then Equation (5.1) shows that every space of directions is homeomorphic to a sphere. Since the link also has spherical links, the inductive assumption and Equation (5.1) show that every $k$-th iterated space of directions ($k\geq 2$) is homeomorphic to a sphere. ∎
@@
Let $X$ be a polyhedral Alexandrov homology $n$-manifold. For any integer $1\leq k\leq n$, let $E_{k}$ be the set of points $p\in X$ such that some $k$-th iterated space of directions at $p$ is not homeomorphic to $S^{n-k}$. If $E_{k}$ is nonempty, then it is an extremal subset of $X$ that is a subcomplex of $X$ only with $k$- or (k1)(k-1)-dimensional maximal simplexes.
@@
The proof only uses the properties of polyhedral spaces and Edwards’ characterization of topological manifolds [9, §8] (cf. Properties (1)–(4) in Section 4). We do not rely on any properties of Alexandrov spaces.
Let $U$ be a neighborhood of $p$ that is isometric to a neighborhood of the vertex of the $\kappa$-cone over $\Sigma_{p}X$. Let $x\in U$ be a point in the direction $\xi\in\Sigma_{p}X$. Then we have
@@
By the inductive assumption and Claim 5.3, for any $p\in E_{k}(X)$, the space of directions $\Sigma_{p}E_{k}(X)$ is an extremal subset of $\Sigma_{p}X$ that is a subcomplex of $\Sigma_{p}X$ only with (k1)(k-1)- or (k2)(k-2)-dimensional maximal simplexes. This implies that $E_{k}(X)$ only have $k$- or (k1)(k-1)-dimensional maximal simplexes. Furthermore, by Lemma 4.2, $E_{k}(X)$ is an extremal subset of $X$. Note that the additional conditions in Lemma 4.2 are satisfied. Indeed, Claim 5.3 implies that $\Sigma_{p}E_{k}(X)$ is nonempty (since $k\geq 2$), and if $\Sigma_{p}E_{k}(X)$ consists of one element (when $k=2$), the same argument as in Step 1 of the proof of Claim 4.3 shows that $\operatorname{diam}\Sigma_{p}E_{k}(X)\leq\pi/2$. This completes the proof. ∎
@@
Let $\Sigma$ be a polyhedral Alexandrov homology manifold with curvature $\geq 1$. Suppose $\Sigma$ contains $k+3$ points with pairwise distance $\pi/2$. Then every $k$-th iterated space of directions is homeomorphic to a sphere.
@@
Let $P$ be a truncated cube obtained by cutting off all the corners of a solid cube by planes. Let $p\in P$ a midpoint of an edge of the original cube. Consider the double $D(P)$ of $P$, i.e., the gluing of two copies of $P$ along the boundary. Then $D(P)$ contains no proper extremal subsets. On the other hand, the space of directions $\Sigma_{p}D(P)$ is isometric to the double of the quadrant of the unit sphere, which contains a 0-dimensional extremal subset. A similar example can be obtained by taking the intersection of the solid cube in $\mathbb{R}^{3}$ with edge length $2$ and the concentric Euclidean ball of radius $\sqrt{2}$ and doubling it. In this case the resulting space has no proper extremal subsets even “locally” around the midpoint of the original edge.
@@
The $xw$-, ywyw-, and $zw$-planes are primitive extremal subsets of $X$. In particular, their union $E$ is an extremal subset of $X$, but not primitive. The projection $\pi(E)$ is an extremal subset of codimension $2$ in $X/\mathbb{Z}$ ([33, 4.1]). Furthermore, $\pi(E)$ is primitive because it can no longer be divided into each plane, due to the effect of the group action. However, $\pi(E)$ does not admit any Alexandrov metric with a lower curvature bound as its intrinsic geodesics branch.